# Inference in Tightly Identified and Large-Scale Sign-Restricted SVARs

### 24 Apr 2026 v1

@@ 1 Introduction
However, the performance of the accept-reject algorithm crucially depends on the size of the admissible set of orthogonal matrices satisfying the restrictions. As the restrictions become tighter, the admissible set shrinks and the probability of accepting a draw declines sharply, rendering the algorithm increasingly inefficient and, in extreme cases, infeasible. This concern is particularly relevant because in recent contributions, strong identifying restrictions that substantially reduce the size of the admissible set are increasingly being imposed [[, see, e.g.,], and the references therein]inoue2026. In the same vein, accept–reject algorithms face challenges in large-scale SVAR models, where the parameter space is high-dimensional and identification often requires imposing a large number of sign restrictions, frequently combined with additional identifying constraints. For example, [chan2025] impose more than one hundred sign restrictions to identify eight shocks in a 35-variable SVAR. Such designs naturally entail a substantial computational burden and highlight the need for inference methods that extend beyond standard accept-reject procedures.
@@
Instead of relying on accept-reject algorithms, we propose imposing inequality restrictions through a reparameterization of the structural model by continuously differentiable mappings. The restrictions are thus imposed by construction, which eliminates the need to discard posterior draws and substantially improves computational efficiency, especially when the admissible set is small. Our approach accommodates a broad range of inequality restrictions commonly used in the SVAR literature. These restrictions, applicable in both small and large models, include shape and ranking restrictions on impulse responses at multiple horizons, as well as bounds on the magnitude of economically relevant elasticities. The framework also allows imposing zero restrictions without requiring importance weighting or other corrections typically needed under accept-reject sampling.
@@
The flexibility of the proposed reparameterization comes at a cost, as the induced posterior distribution of the parameters does not belong to a standard family, which precludes direct sampling. Therefore, posterior inference must rely on Markov chain Monte Carlo (MCMC) methods, and we recommend the No-U-Turn Sampler (NUTS) of [hoffman2014], a modern variant of Hamiltonian Monte Carlo (HMC) that has only rarely been employed in the SVAR literature. It exploits gradients of the log-posterior and the smooth posterior geometry induced by the reparameterization to explore the posterior effectively. However, gradient-based sampling is computationally demanding, as each iteration requires repeated evaluations of the log-posterior and its gradient. Ensuring fast computation of the posterior density is therefore crucial, and we show how it can be evaluated efficiently under the proposed reparameterization to facilitate inference even in high-dimensional parameter spaces. For numerical work, we use the software package Stan [stanreference], containing an implementation of the NUTS with several desirable features, including adaptive tuning, to facilitate estimation.
@@
Recently, [kitagawa2025] have argued that standard MCMC algorithms do not perform well in Bayesian inference on set-identified models because set identification induces flat posterior regions along observationally equivalent parameter manifolds, which can severely impair the mixing of these algorithms. However, while they show that random-walk Metropolis performs poorly in this setting, with mixing deteriorating sharply as the size of the identified set increases, those results do not necessarily carry over to modern, gradient-based samplers with adaptive tuning. Indeed, using the simulation design considered in [kitagawa2025], we show that the No-U-Turn Sampler as implemented in Stan is able to effectively explore high-dimensional, set-identified posteriors with flat directions, producing well-mixed chains even when the identified set is large. [^1] The simulation exercise does not actually involve an SVAR model, but in tightly identified SVARs the identification problem is less severe, which suggests that Hamiltonian Monte Carlo is particularly well suited for posterior exploration in set-identified SVAR models, and this conclusion is also supported by our empirical results.
@@
We are not the first to depart from standard accept–reject algorithms in sign-identified SVAR inference, as already [baumeister2015, baumeister2019] propose to use an MCMC sampler to draw the structural parameters directly. Their approach relies on conjugate prior structures, which produce closed-form conditional posteriors for all parameters except those of the contemporaneous structural relations. The latter are first drawn from their marginal posterior, after which inference for the remaining parameters proceeds using their conditional posteriors. This simplifies computation, but comes at the cost of reduced prior flexibility. Moreover, without substantial computational refinements, their algorithm can become computationally demanding in large VAR systems. More recently [hou2024] derives the conditional posteriors for the structural parameters under linear equality and inequality restrictions that can be used to build a Gibbs sampler for an SVAR model. His approach has the limitation that for longer horizons of the IRFs, an accept-reject step is needed, which is not the case with our approach.
@@
[bruns2023] develop an importance sampler for sign-restricted Bayesian SVARs that uses proposal draws based on a conjugate reduced-form posterior. Their approach improves sampling efficiency and enables estimation beyond the standard conjugate prior setup. However, importance sampling methods can perform poorly in high-dimensional settings, and, as the authors note, the approach is not intended for large VAR systems. [read2025] employ slice sampling for posterior simulation and replace binary sign restrictions with a continuous relaxation. Also their algorithm involves an importance sampling step, which is used to obtain draws from the correct target distribution. Furthermore, although their method yields independent posterior draws, it is formulated under a conditionally uniform prior in the way described in [uhlig2017], which is not invariant to the set of imposed restrictions, as shown by [arias2025elliptical]. Our approach instead uses a uniform prior over orthogonal matrices and therefore guarantees invariance to the restrictions (see [arias2025uniform]).
@@
The inference methods recently put forward by [chan2025] and [arias2025elliptical] are particularly useful in tightly identified and large-scale SVARs. [chan2025] exploits the fact that a uniform distribution over the orthogonal matrices is preserved under reordering the columns or flipping their signs, and thus yields multiple parameter draws at a lower computational cost in an accept-reject algorithm. [arias2025elliptical] propose embedding elliptical slice sampling within a Gibbs sampler, which allows draws of the orthogonal matrix to remain within the identified set by construction. Both approaches deliver substantial computational gains in tightly identified and large-scale SVARs, while maintaining the conventional reduced-form prior structure. However, because of reliance on an accept–reject framework, the approach of [chan2025] is not able to achieve the same level of performance as that of [arias2025elliptical].
@@
As the method of [arias2025elliptical] represents the current state of the art in inference in tightly identified and large-scale SVAR models, we focus on benchmarking our approach against theirs in Section 5. In particular, we revisit their empirical applications to an oil market SVAR identified using sign and elasticity restrictions, as well as a 35-variable SVAR model in which ten shocks are identified by sign and ranking restrictions. In both settings, our HMC approach exhibits faster convergence, as measured by the $\hat{R}$ statistic of [vehtari2021]. This is likely to follow from the fact that their Gibbs sampler generates Markov chains with extremely persistent serial correlation, which remains sizeable even at lags in the hundreds, whereas autocorrelations in our sampler decay comparatively quickly.
@@ 2 Model
The usual objects of interest in SVAR analysis are the structural impulse response functions (IRF), which trace out the responses of observables at different time horizons $t+i$, $i=0,1,2,\dots$, for a shock occurring at time $t$. The structural IRFs of the SVAR model are obtained recursively as $\Psi_{0}=A_{0}^{-1}=B$, and
@@ 3 Identifying Restrictions
To identify the structural parameters of an SVAR model, the researcher must use prior information, typically in the form of identifying restrictions. Currently, imposing sign restrictions is probably the most widely used method of identification. In the implementation of [uhlig2005, rubioramirez2010], candidate draws of the impact matrix are generated from the product $PQ$, where $P$ is the Cholesky factor of the posterior covariance matrix $\Sigma$ of $u_{t}$ and $Q$ is an orthogonal rotation matrix drawn from its prior distribution, and only draws satisfying the sign restrictions are kept. However, it is well known that sign restrictions do not point-identify the model, and hence, the prior is not updated within the identified set in Bayesian estimation [[, see, e.g.,]]baumeister2015. This problem can be avoided by sufficiently tight identification of the orthogonal rotation matrix $Q$, as in that case most of the posterior uncertainty comes from uncertainty about the reduced-form parameters [inoue2020].
@@
Instead of sampling $Q$, the method proposed in this paper works with the impact matrix $B$ directly, which facilitates imposing sign restrictions on the parameters of interest using continuously differentiable mappings to transform them into an unconstrained parameter space. Thus, there is no need to perform any rejection sampling. This approach works well even in very tightly identified models, as regardless of the size of the identified set, the sampler always runs on an unrestricted parameter space. Our method allows for sign restrictions on the impulse responses also beyond the impact effect, as long as the horizon to be restricted does not exceed the VAR lag length $p$. We believe that this limitation on the horizon is unlikely to be important in most practical applications, especially in light of recent work by [olea2025], which recommends using relatively long lag lengths.
Besides sign restrictions, our approach facilitates imposing other kinds of identifying restrictions. In contrast to [arias2018], exclusion restrictions can be imposed directly during sampling without any need for an importance sampling step. Also, the effect of a given shock can be constrained to increase over a specified horizon, or a shock can be required to have the largest impact effect on a given variable. The exact method of placing the latter kinds of restrictions will be presented in Section 4.2. Finally, sign restrictions can be combined with other kinds of restrictions, including restrictions on elasticities, as in [kilian2012], who impose them to avoid implausibly large estimates of the short-run elasticity of oil supply under sign identification alone. More generally, we can impose ranking restrictions requiring a shock to have a greater or smaller effect on a given variable than on another variable (see [amirahmadi2021]). The use of these kinds of restrictions is illustrated in Sections 5.1 and 5.2, respectively.
@@ 4.1 Prior and Posterior
Our aim is to sample from the structural parameters of the model, which requires a transformation of the posterior density of the orthogonal reduced-form parameters into the implied density over the structural parameters of interest. Related transformations from the orthogonal reduced-form parameters to the structural objects have been derived in earlier work. For instance, [arias2018] characterize the induced distribution of the structural parameters in the A-model, referring to it as the normal–generalized normal distribution, while [inoue2013] derive the distribution of the first $p+1$ structural impulse response matrices. [inoue2022] provide a further generalization by allowing the number of IRF matrices to differ from $p+1$.
While we build on [inoue2013, inoue2022], we adapt their results to our setting in which the posterior is parameterized in terms of the first $k$ impulse response matrices $\Psi_{0},\Psi_{1},\dots,\Psi_{k}$ (on which the identifying restrictions are imposed) and the remaining VAR coefficient matrices $(\tilde{A}_{k+1},\dots,\tilde{A}_{p})$. As our interest is in estimating a full SVAR model, we keep the VAR matrices in the density, as opposed to [inoue2022] who only consider the IRF parameters. This mixed parameterization also improves the efficiency of posterior simulation because the VAR coefficients tend to exhibit weaker posterior dependence than impulse response parameters. In addition, by exploiting the properties of the $LQ$ decomposition, we obtain a simpler expression for the associated Jacobian determinant than [inoue2013]. Our expression does not include the orthogonal matrix $Q$, which also saves on calculations as we do not need to evaluate a determinant of another large matrix.
@@
The proof of Proposition 1 does not rely on the distribution over the reduced-form parameters being N-IW, and thus any other prior or error distribution can be used as well. However, we state this proposition using the N-IW density for compatibility with much of the existing literature, and because it allows fast estimation of large SVARs, as will be demonstrated in Section 5.2
As our goal is to incorporate both inequality and zero restrictions, we also need a result that gives us the density function under these restrictions. The following proposition is a direct result of standard facts, but we state it here for clarity.
@@
Suppose the random vector $\theta$ has a continuous density over $\mathbb{R}^{n_{1}+n_{2}+n_{3}}$, and partition it as $\theta=(\theta_{1},\theta_{2},\theta_{3})$, with the dimensions of $\theta_{i}$ being $n_{i}$. Let $R\subset\mathbb{R}^{n_{1}}$ be a set of positive probability. We use $f$ to denote a density function. Suppose that $c\in\mathbb{R}^{n_{2}}$ is a vector of constants, and suppose that $\int_{R}\int_{\mathbb{R}^{n_{3}}}f(\theta_{1},c,\theta_{3})\operatorname{d}\!\theta_{3}\operatorname{d}\!\theta_{1}>0$. Then the conditional density satisfies
@@
In the context of our SVAR model, the density $f$ is the density of the structural parameters in (7). The set-identifying restrictions (like sign and ranking restrictions) and exclusions restrictions are imposed on $\theta_{1}$ and $\theta_{2}$, respectively, while $\theta_{3}$ collects all free parameters of the model. The key point of Proposition 2 is that evaluation of the conditional density under the exclusion restrictions requires no normalizing constants. Therefore, in the case of a Metropolis-style algorithm like HMC, it is sufficient to evaluate the unconstrained density on the subspace satisfying the restrictions. It is important to point out that Proposition 2 is more general than our method affords. In particular, we cannot restrict $\theta_{1}$ in a totally arbitrary way, as discussed at the end of the previous section. We also confine equality restrictions to zero restrictions to preculde the possibility of disjoint posteriors [[, see]]kitagawa2025.

### 21 May 2026 v2

@@ 1 Introduction
However, the performance of the accept-reject algorithm crucially depends on the size of the admissible set of orthogonal matrices satisfying the restrictions. As the restrictions become tighter, the admissible set shrinks and the probability of accepting a draw declines sharply, rendering the algorithm increasingly inefficient and, in extreme cases, infeasible. This concern is particularly relevant because strong identifying restrictions that substantially reduce the size of the admissible set are increasingly being imposed in empirical applications [[, see, e.g.,], and the references therein]inoue2026. In the same vein, accept–reject algorithms face challenges in large-scale SVAR models, where the parameter space is high-dimensional and identification often requires imposing a large number of sign restrictions, frequently combined with additional identifying constraints. For example, [chan2025] impose more than one hundred sign restrictions to identify eight shocks in a 35-variable SVAR. Such designs naturally entail a substantial computational burden and highlight the need for inference methods that extend beyond standard accept-reject procedures.
@@
Instead of relying on accept-reject algorithms, we propose imposing inequality restrictions through a reparameterization of the structural model by continuously differentiable mappings. The restrictions are thus directly incorporated into the model, which eliminates the need to discard posterior draws and substantially improves computational efficiency, especially when the admissible set is small. Our approach accommodates a broad range of inequality restrictions commonly used in the SVAR literature. These restrictions, applicable in both small and large models, include shape and ranking restrictions on impulse responses at multiple horizons, as well as bounds on the magnitude of economically relevant elasticities. The framework also allows imposing zero restrictions without requiring importance weighting or other corrections typically needed under accept-reject sampling.
@@
The flexibility of the proposed reparameterization comes at a cost, as the induced posterior distribution of the parameters does not belong to a standard family, which precludes direct sampling. Therefore, posterior inference must rely on Markov chain Monte Carlo (MCMC) methods, and we recommend the No-U-Turn Sampler (NUTS) of [hoffman2014], a modern variant of Hamiltonian Monte Carlo (HMC) that has only rarely been employed in the SVAR literature. It exploits gradients of the log-posterior and the smooth posterior geometry induced by the reparameterization to explore the posterior effectively. However, gradient-based sampling is computationally demanding, as each iteration requires repeated evaluations of the log-posterior and its gradient. Ensuring fast computation of the posterior density is therefore crucial, and we show how it can be evaluated efficiently under the proposed reparameterization to facilitate inference even in high-dimensional parameter spaces. For numerical work, in this paper, we use the software package Stan [stanreference], containing an implementation of the NUTS with several desirable features, including adaptive tuning, to facilitate estimation.
@@
Recently, [kitagawa2025] have argued that standard MCMC algorithms do not perform well in Bayesian inference on set-identified models because set identification induces flat posterior regions along observationally equivalent parameter manifolds, which can severely impair mixing in these algorithms. However, while they show that random-walk Metropolis performs poorly in this setting, with mixing deteriorating sharply as the size of the identified set increases, those results do not necessarily carry over to modern, gradient-based samplers with adaptive tuning. Indeed, using the simulation design considered in [kitagawa2025], we show that the No-U-Turn Sampler as implemented in Stan is able to effectively explore high-dimensional, set-identified posteriors with flat directions, producing well-mixed chains even when the identified set is large. [^1] The simulation exercise does not actually involve an SVAR model, but in tightly identified SVARs the identification problem is less severe, which suggests that Hamiltonian Monte Carlo is particularly well suited for posterior exploration in set-identified SVAR models, and this conclusion is also supported by our empirical results.
@@
We are not the first to depart from standard accept–reject algorithms in sign-identified SVAR inference, as already [baumeister2015, baumeister2019] propose to use an MCMC sampler to draw the structural parameters directly. Their approach relies on conjugate prior structures, which produce closed-form conditional posteriors for all parameters except those of the contemporaneous structural relations. The latter are first drawn from their marginal posterior, after which inference for the remaining parameters proceeds using their conditional posteriors. This simplifies computation, but comes at the cost of reduced prior flexibility. Moreover, without substantial computational refinements, their algorithm can become computationally demanding in large VAR systems. More recently, [hou2024] derives the conditional posteriors for the structural parameters under linear equality and inequality restrictions that can be used to build a Gibbs sampler for an SVAR model. His approach has the limitation that for longer horizons of the IRFs, an accept-reject step is needed, which is not the case with our approach.
@@
[bruns2023] develop an importance sampler for sign-restricted Bayesian SVARs that uses proposal draws based on a conjugate reduced-form posterior. Their approach improves sampling efficiency and enables estimation beyond the standard conjugate prior setup. However, importance sampling methods can perform poorly in high-dimensional settings, and, as the authors note, the approach is not intended for large VAR systems. [read2025] employ slice sampling for posterior simulation and replace binary sign restrictions with a continuous relaxation. Also their algorithm involves an importance sampling step, which is used to obtain draws from the correct target distribution. Furthermore, although their method yields independent posterior draws, it is formulated under a conditionally uniform prior in the way described in [uhlig2017], which is not invariant to the set of imposed restrictions, as shown by [arias2025elliptical]. Instead, our approach uses a uniform prior over orthogonal matrices and therefore guarantees invariance to the restrictions (see [arias2025uniform]).
@@
The inference methods recently put forward by [chan2025] and [arias2025elliptical] are particularly useful in tightly identified and large-scale SVARs. [chan2025] exploits the fact that a uniform distribution over the orthogonal matrices is preserved under reordering the columns or flipping their signs and thus yields multiple parameter draws at a lower computational cost in an accept-reject algorithm. [arias2025elliptical], in turn, propose embedding elliptical slice sampling within a Gibbs sampler, which allows draws of the orthogonal matrix to remain within the identified set by construction. Both approaches deliver substantial computational gains in tightly identified and large-scale SVARs, while maintaining the conventional reduced-form prior structure. However, because of reliance on an accept–reject framework, the approach of [chan2025] is not able to achieve the same level of performance as that of [arias2025elliptical].
@@
As the method of [arias2025elliptical] represents the current state of the art in inference in tightly identified and large-scale SVAR models, we focus on benchmarking our approach against theirs in Section 5. In particular, we revisit their empirical applications to an oil market SVAR identified by sign and elasticity restrictions, as well as a 35-variable SVAR model in which ten shocks are identified by sign and ranking restrictions. In both settings, our HMC approach exhibits faster convergence, as measured by the $\hat{R}$ statistic of [vehtari2021]. While differences are small in the oil market SVAR, with our method only doubling the speed of convergence, they are substantially greater in the 35-variable model. In that setting, our approach converges to the stationary distribution approximately five times faster than the method of [arias2025elliptical]. Measured by the effective sample size (ESS), the number of independent draws equivalent to the correlated MCMC samples, it also attains substantially higher sampling efficiency, with ESS up to an order of magnitude greater than those of the elliptical slice sampler. These results are likely to follow from the fact that the Gibbs sampler of [arias2025elliptical] generates Markov chains with extremely persistent serial correlation, which remains sizeable even at lags in the hundreds, whereas autocorrelations in our sampler decay comparatively quickly.
@@
Finally, it is worth noting that our paper is also related to [plagborgmoeller2019], who estimates vector moving average (VMA) models using the NUTS to sample directly from a moving-average representation. In our SVAR setting, we adopt a mixed parameterization: the moving-average coefficient matrices subject to the identifying restrictions are sampled directly, whereas the remaining dynamics are parameterized in terms of reduced-form VAR coefficients. In our experience, this parametrization is well suited to the NUTS, with chains mixing well and converging rapidly even under diffuse priors. Furthermore, this parameterization allows for the usage of a variety of priors designed for VAR models. By contrast, posterior simulation in unrestricted VMA models is computationally challenging, and satisfactory mixing and convergence may be difficult to achieve without strongly informative priors, which can be restrictive in empirical applications.
@@ 2 Model
The usual objects of interest in SVAR analysis are the structural impulse response functions (IRF), which trace out the responses of observables at different time horizons $t+i$, $i=0,1,2,\dots$, for a shock occurring at time $t$. The structural IRFs are obtained recursively as $\Psi_{0}=A_{0}^{-1}=B$, and
@@ 3 Identifying Restrictions
To identify the structural parameters of an SVAR model, the researcher must use prior information, typically in the form of identifying restrictions. Currently, sign restrictions are probably the most widely used type of such restrictions. In the implementation of [uhlig2005, rubioramirez2010], candidate draws of the impact matrix are generated from the product $PQ$, where $P$ is the Cholesky factor of the posterior covariance matrix $\Sigma$ of $u_{t}$ and $Q$ is an orthogonal rotation matrix drawn from its prior distribution, and only draws satisfying the sign restrictions are retained. However, it is well known that sign restrictions do not point-identify the model, and hence, the prior is not updated within the identified set in Bayesian estimation [[, see, e.g.,]]baumeister2015. This problem can be avoided by sufficiently tight identification of the orthogonal rotation matrix $Q$, because in that case posterior uncertainty is mostly attributed to uncertainty about the reduced-form parameters [inoue2020].
@@
Instead of sampling $Q$, the method proposed in this paper works with the impact matrix $B$ directly, which facilitates imposing sign restrictions on the parameters of interest using continuously differentiable mappings to transform them into an unconstrained parameter space. Thus, there is no need to perform any rejection sampling. This approach works well even in very tightly identified models, as regardless of the size of the identified set, the sampler always runs on an unrestricted parameter space. Our method allows for sign restrictions on the impulse responses also beyond the impact effect, as long as the horizon to be restricted does not exceed the VAR lag length $p$. We believe that this limitation on the horizon is unlikely to be important in most empirical applications, especially in light of the recent work by [olea2025] recommending relatively long lag lengths.
Besides sign restrictions, our approach facilitates imposing other kinds of identifying restrictions. In contrast to [arias2018], exclusion restrictions can be imposed without an importance sampling step, thus avoiding computation of volume elements. Also, the effect of a given shock can be constrained to increase over a specified horizon, or a shock can be required to have the largest impact effect on a given variable. The exact method of placing the latter kinds of restrictions will be presented in Section 4.2. Finally, sign restrictions can be combined with other kinds of restrictions, including restrictions on elasticities, as in [kilian2012], who impose them to avoid implausibly large estimates of the short-run elasticity of oil supply under sign identification alone. More generally, we can impose ranking restrictions requiring a shock to have a greater or smaller effect on a given variable than on another variable (see [amirahmadi2021]). The use of these kinds of restrictions is illustrated in Sections 5.1 and 5.2, respectively.
@@ 4.1 Prior and Posterior
Our aim is to sample from the structural parameters of the model, which requires a transformation of the posterior density of the orthogonal reduced-form parameters into the implied density over the structural parameters of interest. Related transformations from the orthogonal reduced-form parameters to the structural parameters have been derived in previous work. For instance, [arias2018] characterize the induced distribution of the structural parameters in the A-model, referring to it as the normal–generalized normal distribution. [inoue2013] derive the distribution of the first $p+1$ structural impulse response matrices, and [inoue2022] provide a further generalization by allowing the number of IRF matrices to differ from $p+1$.
While we build on [inoue2013, inoue2022], we adapt their results to our setting in which the posterior is parameterized in terms of the first $k$ impulse response matrices $\Psi_{0},\Psi_{1},\dots,\Psi_{k}$ (on which the identifying restrictions are imposed) and the remaining VAR coefficient matrices $(\tilde{A}_{k+1},\dots,\tilde{A}_{p})$. As our interest is in estimating a full SVAR model, we keep the VAR matrices in the density, as opposed to [inoue2022] who only consider the IRF parameters. This mixed parameterization also improves the efficiency of posterior simulation because the VAR coefficients tend to exhibit weaker posterior dependence than impulse response parameters. In addition, by exploiting the properties of the $LQ$ decomposition, we obtain a simpler expression for the associated Jacobian determinant than [inoue2013]. It does not include the orthogonal matrix $Q$, which also saves on calculations as there is no need to evaluate a determinant of another large matrix.
@@
The proof of Proposition 1 does not rely on the distribution over the reduced-form parameters being N-IW, and thus any other prior or error distribution can be specified as well. For example, the asymmetric conjugate prior of [chan2022] could be used with an appropriate Jacobian adjustment. However, we state this proposition using the N-IW density for compatibility with much of the existing literature, and because it allows fast estimation of large SVARs, as will be demonstrated in Section 5.2
Our goal is to incorporate both inequality and zero restrictions, and thus we also need a result that gives us the density function under these restrictions. The following proposition is a direct result of standard facts, but we state it here for clarity.
@@
Suppose the random vector $\theta$ has a continuous density over $\mathbb{R}^{n_{1}+n_{2}+n_{3}}$, and partition it as $\theta=(\theta_{1},\theta_{2},\theta_{3})$, with the dimensions of $\theta_{i}$ being $n_{i}$. Let $R\subset\mathbb{R}^{n_{1}}$ be a set of positive probability. Suppose that $c\in\mathbb{R}^{n_{2}}$ is a vector of constants, and $\int_{R}\int_{\mathbb{R}^{n_{3}}}f(\theta_{1},c,\theta_{3})\operatorname{d}\!\theta_{3}\operatorname{d}\!\theta_{1}>0$. Then the conditional density of $\theta$ satisfies
@@
In the context of our SVAR model, the density $f$ is the density of the structural parameters $\theta$ in (7). The set-identifying restrictions (like sign and ranking restrictions) and exclusion restrictions are imposed on $\theta_{1}$ and $\theta_{2}$, respectively, while $\theta_{3}$ collects all free parameters of the model. The key point of Proposition 2 is that evaluation of the conditional density under the exclusion restrictions requires no normalizing constants. Therefore, in the case of a Metropolis-style algorithm like HMC, it is sufficient to evaluate the unconstrained density on the subspace satisfying the restrictions. It is important to point out that Proposition 2 is more general than our method affords. In particular, we cannot restrict $\theta_{1}$ in a totally arbitrary way, as discussed at the end of Section 3. We also confine equality restrictions to zero restrictions to preculde the possibility of disjoint posteriors [[, see]]kitagawa2025.